OpenStudy (anonymous):

How would you solve y=- (0)+2?

OpenStudy (jdoe0001):

solve?

OpenStudy (anonymous):

y = +2

OpenStudy (anonymous):

how did you get y=+2?

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

you can ignore ZERO

OpenStudy (anonymous):

it has no value

OpenStudy (anonymous):

And where did the - go?

OpenStudy (jdoe0001):

\(\bf y=- (0)+2\implies y=-1\cdot (0)+2\implies y=0+2\implies y=2\)

OpenStudy (anonymous):

Thanks a ton :)

OpenStudy (jdoe0001):

yw